# Solutions for $k$-generalized Fibonacci numbers using Fuss-Catalan numbers

### v3

@@ 6.6 Perrin sequence
The Perrin and Padovan sequences satisfy the same recurrence, but with different initial values. We write ‘$Q(n)$’ to avoid confusion with the Padovan sequence. Then $Q(0)=3$, $Q(1)=0$ and $Q(2)=2$ and for $n\geq 3$ the recurrence is
$Q(n)=Q(n-2)+Q(n-3)\,.$ (6.10)
Recall the sum in eq. (6.5) is normalized to yield $F(k-1)=F(2)=1$. Setting $F(0)=1$ and $F(1)=0$ yields
$Q(n)=2F(n)+P(n-5)=2P(n-2)+P(n-5)\,.$ (6.11)
Here $F(n)$ is given by eq. (6.9) and recall $P(n)$ is a Padovan number. For $n<5$, one must extend the Padovan sequence to negative arguments. The first few terms are $3,0,2,3,2,5,5,7,10,12,17,22,\dots$.
7 Generating function
The generating function for the $k$-generalized Fibonacci numbers is
$G(z)=\frac{z^{k-1}}{1-z-\dots-z^{k}}\,.$ (7.1)
Then $F_{n}(n)$ is the coefficient of $z^{n}$ in the series expansion of $G(z)$. Process the generating function as follows
$\begin{split}G(z)&=\frac{z^{k-1}}{1-z-\dots-z^{k}}\\ &=\frac{z^{k-1}}{1-z(1+\dots+z^{k-1})}\\ &=\frac{z^{k-1}}{1-z(1-z^{k})/(1-z)}\\ &=\frac{(1-z)z^{k-1}}{1-2z+z^{k+1}}\,.\end{split}$ (7.2)
Define $T=1/(1-2z+z^{k+1})$ and expand it in a series as follows
$\begin{split}T&=\frac{1}{1-2z+z^{k+1}}\\ &=\frac{1}{1-2z}\frac{1}{1+z^{k+1}/(1-2z)}\\ &=\frac{1}{1-2z}\sum_{s=0}^{\infty}(-1)^{s}\frac{z^{s(k+1)}}{(1-2z)^{s}}\\ &=\sum_{s=0}^{\infty}(-1)^{s}\frac{z^{s(k+1)}}{(1-2z)^{s+1}}\\ &=\sum_{s=0}^{\infty}(-1)^{s}z^{s(k+1)}\sum_{u=0}^{\infty}\binom{s+u}{s}(2z)^{u}\\ &=\sum_{j=0}^{\infty}z^{j}\sum_{s=0}^{\lfloor s/(k+1)\rfloor}(-1)^{s}2^{j-s(k+1)}\binom{j-sk}{s}\qquad(j=s(k+1)+u)\,.\end{split}$ (7.3)
Then $G(z)$ consists of a sum of two series
$\begin{split}G(z)&=z^{k-1}\sum_{j=0}^{\infty}z^{j}\sum_{s=0}^{\lfloor j/(k+1)\rfloor}(-1)^{s}2^{j-s(k+1)}\binom{j-sk}{s}\\ &\quad-z^{k}\sum_{j=0}^{\infty}z^{j}\sum_{s=0}^{\lfloor j/(k+1)\rfloor}(-1)^{s}2^{j-s(k+1)}\binom{j-sk}{s}\,.\end{split}$ (7.4)
The coefficient of $z^{n}$ is $F_{k}(n)$:
$\begin{split}F_{k}(n)&=\sum_{s=0}^{\lfloor(n-k+1)/(k+1)\rfloor}(-1)^{s}2^{n-k+1-s(k+1)}\binom{n-k(1+s)+1}{s}\\ &\quad-\sum_{s=0}^{\lfloor(n-k)/(k+1)\rfloor}(-1)^{s}2^{n-k-s(k+1)}\binom{n-k(1+s)}{s}\,.\end{split}$ (7.5)
We can set both upper limits equal to those in the first sum. (This may introduce an extraneous term in the second sum, but if so, it equals zero.) The final result is more elegant.
$F_{k}(n)=\sum_{s=0}^{\lfloor(n-k+1)/(k+1)\rfloor}(-1)^{s}2^{n-k-s(k+1)}\biggl{[}2\binom{n-k(1+s)+1}{s}-\binom{n-k(1+s)}{s}\biggr{]}\,.$ (7.6)
Observe that the sum is not nested, hence the computational complexity to evaluate $F_{k}(n)$ does not increase as the value of $k$ increases.
Remark 7.1.
A similar expression was obtained by Ferguson [12], where the series expansion of $T$ in eq. (7.3) was derived differently. See also Kessler and Schiff ([13] eq. (6)), which generalized Ferguson’s work beyond $k$-generalized Fibonacci numbers.
8 Basis sequences
Let us return to the Vandermonde system of equations in eq. (3.1). The $k$-generalized Fibonacci numbers are the solution for only one set of initial conditions, viz. $F_{k}(0)=\dots=F_{k}(k-2)=0$ and $F_{k}(k-1)=1$. Let us derive a set of $k$ ‘basis sequences’ $B_{km}(n)$, $m=0,\dots,k-1$ as follows. First, for consistency of notation, we reexpress the recurrence eq. (1.1) as
$B_{k}(n)=B_{k}(n-1)+\dots+B_{k}(n-k)\,.$ (8.1)
The basis sequence $B_{km}(n)$ is the solution of eq. (8.1) for the the initial values $B_{km}(n)=\delta_{mn}$ for $10\leq m,n\leq k-$ 1. (The $k$-generalized Fibonacci numbers are the last member of the basis, viz. $m=k-1$, so $F_{k}(n)=B_{k,k-1}(n)$.) Then, for arbitrary initial values $(\gamma_{0},\dots,\gamma_{k-1})$ for $n=0,\dots,k-1$, the solution of the recurrence eq. (8.1) is
$B_{k}(n)=\sum_{m=0}^{k-1}\gamma_{m}B_{km}(n)\,.$ (8.2)
To proceed, we extend the Vandermonde system of equations as follows (attaching row and column indices to the coefficients $c$)
$\begin{split}\begin{pmatrix}1&1&\dots&1\\ \zeta_{0}&\zeta_{1}&\dots&\zeta_{k-1}\\ \zeta_{0}^{2}&\zeta_{1}^{2}&\dots&\zeta_{k-1}^{2}\\ \vdots&&&\vdots\\ \zeta_{0}^{k-1}&\zeta_{1}^{k-1}&\dots&\zeta_{k-1}^{k-1}\end{pmatrix}\begin{pmatrix}c_{00}&c_{01}&\dots&c_{0,k-1}\\ c_{10}&c_{11}&\dots&c_{1,k-1}\\ c_{20}&c_{21}&\dots&c_{2,k-1}\\ \vdots&&&\vdots\\ c_{k-1,0}&c_{k-1,1}&\dots&c_{k-1,k-1}\end{pmatrix}=\begin{pmatrix}1&0&\dots&0\\ 0&1&\dots&0\\ \vdots&\ddots&&\vdots\\ \vdots&&\ddots&\vdots\\ 0&0&\dots&1\end{pmatrix}\,.\end{split}$ (8.3)
The basis sequence $B_{km}(n)$ is given by
$B_{km}(n)=\sum_{\ell=0}^{k-1}c_{\ell m}\zeta_{\ell}^{n}\,.$ (8.4)
To determine $c_{\ell m}$, we invert the Vandermonde matrix in eq. (8.3). The inverse is the matrix whose rows are the coefficients of the Lagrange polynomials $L_{\ell}(x)$, $\ell=0,\dots,k-1$. The solution is $c_{\ell m}=L_{\ell m}$, which is the coefficient of $x^{m}$ in $L_{\ell}(x)$. Recall from eq. (5.1) the Lagrange polynomial $L_{\ell}(x)$ is
$L_{\ell}(x)=\frac{(x-\zeta_{0})\dots(x-\zeta_{\ell-1})(x-\zeta_{\ell+1})\dots(x-\zeta_{k-1})}{(\zeta_{\ell}-\zeta_{0})\dots(\zeta_{\ell}-\zeta_{\ell-1})(\zeta_{\ell}-\zeta_{\ell+1})\dots(\zeta_{\ell}-\zeta_{k-1})}\,.$ (8.5)
We denote the numerator by $N_{\ell}(x)$ and write $N_{\ell}(x)=\sum_{m=0}^{k-1}N_{\ell m}x^{m}$. The coefficients $N_{\ell m}$ have the following structure
  1. 1.
    First, $N_{\ell,k-1}$ equals unity.
  2. 2.
    Next, for $j=1,\dots,k-1$, $N_{\ell,k-1-j}=(-1)^{j}\times$ (sum of the products of $j$-tuples of roots excluding $\zeta_{\ell}$).
  3. 3.
    Hence $N_{\ell,k-2}$ equals the negative of the (sum of the roots excluding $\zeta_{\ell}$) and $N_{\ell,k-3}$ equals the sum of the (products of pairs of roots excluding $\zeta_{\ell}$), etc.
The coefficients $N_{\ell m}$ can be obtained from the auxiliary polynomial $A(x)$ as follows. First, define the sums of products of $j$-tuples of roots as follows $S_{j}=\sum_{0\leq i_{1}<\dots<i_{j}\leq k-1}\zeta_{i_{1}}\dots\zeta_{i_{j}}$, $j=1,\dots,k-1$. Recall $A(x)=x^{k}-x^{k-1}-\dots-1$ and by definition $A(x)=(x-\zeta_{0})\dots(x-\zeta_{k-1})$. From the theory of polynomials, observe that $S_{j}=(-1)^{j-1}$. Then some careful bookkeeping yields
$N_{\ell,k-1}$ $=1\,,$ (8.6a)
$N_{\ell,k-2}$ $=\zeta_{\ell}N_{\ell,k-1}-1$ $=$ $\;\zeta_{\ell}-1\,,$ (8.6b)
$N_{\ell,k-3}$ $=\zeta_{\ell}N_{\ell,k-2}-1$ $=$ $\;\zeta_{\ell}^{2}-\zeta_{\ell}-1\,.$ (8.6c)
In general, $N_{\ell,k-1}=1$ and $N_{\ell,m}=\zeta_{\ell}N_{\ell,m+1}-1$ and we work recursively to deduce, for $m=k-2,k-3,\dots,0$,
$N_{\ell,m}=\zeta_{\ell}^{k-m-1}-\zeta_{\ell}^{k-m-2}-\dots-1\,.$ (8.7)
The expression for the denominator $D_{\ell}$ of the Lagrange polynomial $L_{\ell}(x)$ was derived in eq. (5.4). Then the coefficient $c_{\ell m}$ is given by
$c_{\ell m}=N_{\ell m}\,\frac{\zeta_{\ell}-1}{(k+1)\zeta_{\ell}-2k}\,\zeta_{\ell}^{1-k}\,.$ (8.8)
Finally, the basis sequence $B_{km}(n)$ is given by ($m=0,\dots,k-1$)
$B_{km}(n)=\sum_{\ell=0}^{k-1}N_{\ell m}\,\frac{\zeta_{\ell}-1}{(k+1)\zeta_{\ell}-2k}\,\zeta_{\ell}^{n-k+1}\,.$ (8.9)
This furnishes a basis (or ‘fundamental solutions’) for the recurrence eq. (8.1).
Note the following. The first and last basis sequences $B_{k0}(n)$ and $B_{k,k-1}(n)$ are related via a shift of the index $n$ as follows
$B_{k0}(n)=B_{k,k-1}(n-1)\,.$ (8.10)
Proof.
First note that $B_{k0}(k)=1$. Then observe that $\{B_{k0}(1),\dots,B_{k0}(k-1),B_{k0}(k)\}=\{0,\dots,0,1\}$. These are exactly the initial values for the sequence $B_{k,k-1}(n)$, but the index $n$ is shifted by unity. Eq. (8.10) follows immediately. ∎
As an example, let us apply the above results to the Fibonacci numbers, i.e. $k=2$. Denoting the Fibonacci numbers by $F_{n}$ and using eq. (8.10) yields $B_{2,0}(n)=F_{n-1}$ and $B_{2,1}(n)=F_{n}$. Then for the initial values $(\gamma_{0},\gamma_{1})$, the general solution is $B_{2}(n)=\gamma_{0}F_{n-1}+\gamma_{1}F_{n}$. This proves (yet again) that every solution of the standard Fibonacci recurrence can be expressed as a linear combination of $F_{n-1}$ and $F_{n}$. The Lucas numbers $L_{n}$ have the initial values $(L_{0},L_{1})=(2,1)$, whence
$L_{n}=2F_{n-1}+F_{n}=2F_{n+1}-F_{n}\,.$ (8.11)
This is the known relation of the Lucas numbers to the Fibonacci numbers.

### v4

@@ 9.6 Perrin sequence
The Perrin and Padovan sequences satisfy the same recurrence, but with different initial values. We write ‘$Q(n)$’ to avoid confusion with the Padovan sequence $P(n)$. The initial values are $Q(0)=3$, $Q(1)=0$ and $Q(2)=2$ and for $n\geq 3$ the recurrence is
Q(n)=Q(n2)+Q(n3).Q(n)=Q(n-2)+Q(n-3)\,. (9.11)
Because the recurrence eq. (9.11) is the same as eq. (9.8), the basis functions are the same as in Sec. 9.5, viz. $W^{P}_{i}(n)$, $i=0,1,2$. Using the given initial values, the solution is
$\begin{split}Q(n)&=3W^{P}_{0}(n)+2W^{P}_{2}(n)\\ &=3W^{P}_{2}(n-1)+2W^{P}_{2}(n)\\ &=3\bigl{(}\,\underbrace{W^{P}_{2}(n-1)+W^{P}_{2}(n)}_{=W^{P}_{2}(n+2)}\,\bigr{)}-W^{P}_{2}(n)\\ &=3W^{P}_{2}(n+2)-W^{P}_{2}(n)\,.\end{split}$ (9.12)
We can express the Perrin numbers as a linear combination of the Padovan numbers. One choice (not unique) is Q(n)=4P(n)+2P(n+1)3P(n+2)Q(n)=4P(n)+2P(n+1)-3P(n+2). The first few values of $Q(n)$ are $3,0,2,3,2,5,5,7,10,12,17,22,\dots$.
Acknowledgement
The author is grateful to David Wolfram for numerous helpful comments on the manuscript and especially for bringing Ref. [13] to his attention. The author is also indebted to the referees for valuable suggestions for improvement, including the connection to Dickson polynomials and the reference to Chen and Louck [2].
Disclosure statement
The author declares there are no relevant financial or non-financial competing interests to report.